Babu's age is three times his son's age.Ten years ago he was five times his son's age.Find their present age.
Answer by galactus(183)

Let x=Babu's age now
y=son's age now
Presently, Babu is 3 times his sons age: x=3y
10 years ago he was 5 times his age: x-10=5(y-10)
3y-10=5y-50
2y=40
y=20
His son his 20, and since his father (Babu) is 3 times that, he is 60.
Check. 10 years ago Babu would have been 50 and his son 10. 50 is 5 times 10.
Check!